OpenStudy (anonymous):

Find the amount if $12 000 is invested at 5.75%/a compounded semi-annually for 10 years

OpenStudy (anonymous):

A = P(1+r/n)^nt A=12000(1+ (0.0575/2) )^(2*10) 21 153.304 P = principal amount r = annual rate of interest t = number of years the amount is deposited or borrowed for. A = amount of money accumulated after n years, including interest. n = number of times the interest is compounded per year